OpenStudy (anonymous):

Write an equation of the line that passes through (1,-4) and is perpendicular to the line: y = 2x + 5. Could anyone explain how to solve this? Thank you.

OpenStudy (texaschic101):

Look at your equation y = 2x + 5. First we need to find the slope. In y = mx + b form, the slope is in the m position. So can you tell me the slope of this equation ?

OpenStudy (anonymous):

The slope of the existing line is 2 and I know that since I'm trying to find the equation of a line that's perpendicular to y = 2x + 5 the slope of the perpendicular line would be -1/2. I don't know much beyond that.

OpenStudy (texaschic101):

cool...you got that much correct. Now their is 2 different ways to continue... We can either use y = mx + b form or we can use y - y1 = m(x - x1). I , myself, prefer the y = mx + b form (slope intercept form) unless specifically asked for the other y - y1 = m(x - x1), point slope form. So we will use y = mx + b we already know we need to use slope -1/2. and our points (1,-4)....x = 1 and y = -4 now we will sub....we are looking for the y intercept (b) y = mx + b -4 = -1/2(1) + b -4 = -1/2 + b -4 + 1/2 = b -8/2 + 1/2 = b -7/2 = b so your perpendicular equation is : y = -1/2x - 7/2 if you have any questions, please ask

OpenStudy (texaschic101):

that is the answer in slope intercept form, but if you need it in standard form, let me know and I will continue

OpenStudy (anonymous):

That's all I need, thanks so much for your help.

OpenStudy (texaschic101):

no problem...glad to help :)